Three numbers between 100 and 150 can each be expressed as the sum of two different squares in two different ways. What is the sum of the three numbers?
Answer by ikleyn(52803)

.

These three numbers are


    125 = 2^2 + 11^2 = 5^2 + 10^2,


    130 = 3^2 + 11^2 = 7^2 + 9^2,


    145 = 1^2 + 12^2 = 8^2 + 9^2.


Their sum is  125 + 130 + 145 = 400.    ANSWER

Solved.